OpenStudy (anonymous):

Look at the following number sequence. 6, 13, 20, 27, . . . Based on the pattern, what are the next two terms?

OpenStudy (anonymous):

well I can give u a hint!

OpenStudy (anonymous):

there positive numbers!

OpenStudy (anonymous):

Each term is 7 more than the previous term. Now, what are the two terms?

OpenStudy (anonymous):

34 and 41?

OpenStudy (anonymous):

+7, +7, +7 6+7 =13 13+7 =20 20+7=27...